OpenStudy (anonymous):

how to solve 3a^2+6a+78=0 using solving quadratic equations by completing the square?

OpenStudy (anonymous):

i would divide by 3 first

OpenStudy (anonymous):

\[a^2+2a+26=0\]

OpenStudy (anonymous):

clear answers will be complex right?

OpenStudy (anonymous):

(a+1)^2+25=0

OpenStudy (anonymous):

if that is not a problem we can go right to the answer by completing the square \[(a^2+2a=-26\] \[(a+1)^2=-26+1=-25\] \[a+1=\pm5i\] \[a=-1\pm5i\]